OpenStudy (anonymous):

x<5 and x> 0 interval notation

OpenStudy (anonymous):

(5,0)?

OpenStudy (lgbasallote):

interval notation...is that the one that looks like [a,b]U(b,a) or something like that?

OpenStudy (anonymous):

okay

OpenStudy (lgbasallote):

okay??? im asking you

OpenStudy (anonymous):

oh idk

OpenStudy (lgbasallote):

i dont know those fancy terms i just know what the answer is based on how they look like lol

Parth (parthkohli):

May I answer?

OpenStudy (anonymous):

okay lol

OpenStudy (anonymous):

and yes parth

OpenStudy (lgbasallote):

perhaps you should start by explaining what interval notation is @ParthKohli lol

OpenStudy (anonymous):

\[(-\infty,5)U(0,+\infty)\]

OpenStudy (lgbasallote):

oh..it was the one i was describing

OpenStudy (anonymous):

okay that's how you do it. Thanks

OpenStudy (anonymous):

idk wether i should have added that or not lol

Parth (parthkohli):

Ok...... If it's written x\(\ge\)5, then it'd look like this [5,\(+\infty\)) square bracket means that five is included.

OpenStudy (lgbasallote):

@kswiss is smart..he didnt know what interval notation was and he can deduce how to do it by seeing an answer...very smart ^_^

OpenStudy (anonymous):

okay

Parth (parthkohli):

Now we know that here it can be 1,2,3,4 but now 5 and 0, so we may have something like this: (0,5) ^^ not including 0 till not including 5.

Parth (parthkohli):

Or another way to write it is [1,4]

OpenStudy (anonymous):

gotcha